# An exploratory study on automatic identification of assumptions in the development of deep learning frameworks

### 08 Jan 2024 v1

@@ 4.1 Research questions
Rationale: Algorithms usually have varying performance in classification problems, leading to different trained models. For example, Naive Bayes models are a group of probabilistic classifiers that have been widely used in text categorization since 1960. The Support Vector Machine algorithm is a commonly used supervised ML algorithm for data classification and regression analysis. To answer RQ1, we chose seven popular ML algorithms to train classification models for the identification of assumptions: Perceptron (Pct), Logistic Regression (LR), Linear Discriminant Analysis (LDA), K-Nearest Neighbors (KNN), Support Vector Machine (SVM), Naive Bayes (NB), and Classification and Regression Trees (CART). These algorithms have been used in similar studies, such as the classification of non-functional requirements Abad2017 and rationale Alkadhi2018 . Li et al. also used these seven ML algorithms to identify assumptions (binary classification) in the Hibernate project Li2019 . Through this RQ, we plan to evaluate the performance of traditional ML models for identifying assumptions on the AssuEval dataset yangpackage2023 .
@@
Rationale: In addition to traditional ML and DL models, large language models (LLMs) are another potential solution for identifying assumptions. In this RQ, we chose one of the most representative LLMs (i.e., ChatGPT) to explore its performance in identifying assumptions. ChatGPT is an advanced AI chatbot developed by OpenAI [^3] https://openai.com/, which can interact in a conversational way chatgptintro2022 . As reported by OpenAI, more than 100 million people used ChatGPT [^4] https://chatgptui.org/features?step=1. ChatGPT can be applied in various fields, such as answering specific questions, writing an essay, writing a song, writing code to implement a function, and code debugging chatgptintro2022 Meyer2023 . Specifically, we chose the free version [^5] https://chat.openai.com/chat of ChatGPT (based on GPT-3.5) as it is more available to the public. The results of this RQ show the performance of ChatGPT in identifying assumptions on the AssuEval dataset.
@@ 4.3 Data collection
The process of data collection is shown in Fig. 7.
(1) Select repositories. TensorFlow is one of the most popular DL frameworks, which is widely used in many DL systems and application domains. The TensorFlow project on GitHub started in 2015, having 156,600 commits [^6] https://github.com/tensorflow/tensorflow (accessed on 2023-11-26), 23,530 PRs [^7] https://github.com/tensorflow/tensorflow/pulls?q=is%3Apr (accessed on 2023-11-26), and 38,344 issues [^8] https://github.com/tensorflow/tensorflow/issues?q=is%3Aissue (accessed on 2023-11-26) till November 2023. Therefore, in this study, we chose the TensorFlow repository on GitHub as the main source to build our dataset. Moreover, Keras is an independent repository on GitHub, which is closely related to TensorFlow. For example, in some versions of TensorFlow, Keras has been integrated inside TensorFlow as a sub-module (i.e., tensorflow.keras). Therefore, in addition to TensorFlow, we also included the Keras repository on GitHub when constructing the AssuEval dataset yangpackage2023 .
(2) Determine the scope of the data. According to similar works (e.g., Zhang2018 Ruan2019 Han2020 ) that collect and analyze the data on GitHub, in this study we focus on commits, PRs, and issues of the selected repositories.
@@ 4.4.1 Construction process of the AssuEval dataset
(3) Conduct a pilot labeling. We conducted a pilot labeling on the TensorFlow and Keras repositories. The criteria were refined iteratively during the pilot study. For example, regarding an assumption that is a condition (e.g., “This obviously works if you can assume 0.0 is not real value in your data”), we decided not to label such data as SCA, because the assumption does not exist. Another example is that if a sentence is a question (e.g., “@yaroslavvb I’m assuming that if I run the probe op in a session together with computation of a model this would return me the peek memory usage, is that correct?”), which is about the validity of an assumption rather than whether to make an assumption, we decided to label such data as SCA. More details of the included or excluded examples can be found in Table 2, Table 3, Table 4, and Table 5.
@@
(5) Label the data from the selected repositories. The first author labeled the data following the criteria using Assumption Miner. NAs are labeled as “0”, PAs are labeled as “1”, and SCAs are labeled as “2”.
@@ 4.4.2 Criteria of data labeling
The SCA concept and the PA concept. As the scope of the study is to analyze the DL framework projects on GitHub, before designing the dataset construction criteria, we first classified assumptions into two types: self-claimed assumptions (SCAs) and potential assumptions (PAs). The rationale of such a classification is that a sentence cannot be identified as an assumption unless it is explicitly mentioned by stakeholders, which are SCAs. There are also some plausible cases that may or may not be assumptions, which we have to ask stakeholders to confirm, which are PAs. Taking into account a sentence in a commit, PR, or issue of a GitHub repository, an SCA is that the sentence includes an assumption and the assumption is explicitly claimed using at least one of the assumption-related terms (i.e., “assumption”, “assumptions”, “assume”, “assumes”, “assumed”, “assuming”, “assumable”, and “assumably”). For example, in this sentence “[tf/xla] fixup numbering of xla parameters used for aliasing previously, the xla argument parameter was incorrectly assumed to correspond to the index in the vector of ‘xlacompiler::argument’”, it includes an SCA of assuming that the xla argument parameter corresponds to the index in the vector of xlacompiler::argument and the developer claimed that it was an invalid assumption. A PA is that the sentence may include an assumption, which is a potential assumption that needs further confirmation from human experts. This definition covers various aspects, such as expectations, future events, possibilities, guesses, opinions, feelings, and suspicions, which can indicate PAs. We provide three examples to further explain the concept of PA. For example, in the sentence “I think the right way to create demo tensorboard instances is to simply run a tensorboard in the cloud, rather than keep maintaining this mocked-out backend.”, it includes a PA regarding thoughts on the right way to create demo tensorboard instances. In another example “The system will not crash under heavy load”, the sentence describes the future state of the system, which is uncertain and includes a PA. The third example “both false and true outputs should be considered independently” does not include assumption-related terms, but the sentence describes an expectation (i.e., “something should be”), which is a PA. After further confirmation by human experts, this PA can be transformed into an SCA or other types of software artifacts. Besides SCAs and PAs (which belong to explicit assumptions), there are also many implicit assumptions in projects (e.g., in stakeholders’ heads or requiring reasoning). Identifying implicit assumptions is much more tricky than identifying SCAs and PAs, since there are no explicit clues in the sentences and stakeholders need to infer the sentences based on the context, which involves assumption reasoning. Identifying implicit assumptions is beyond the scope of this study and we treat it as future work. Therefore, for a sentence that is neither an SCA nor a PA, we labeled it as NA.
@@
E1: Rename the pass to merge assuming ops.
@@ 4.5.2 Model construction and training
For answering RQ1, we used scikit-learn to implement and train the seven ML models. The configurations of each model are shown in Table 6.
@@
For answering RQ2, we used the TensorFlow framework to train the ALBERT model. The configurations of the ALBERT model are shown in Table 7. The optimizer is Adam, the training batch size is 64, and the maximum sequence length of each sentence is 128. We trained the model to 10,000 epochs with an initial learning rate of 2e-5.
@@
To answer RQ3, we used the pre-trained ChatGPT model without retraining or further fine-tuning. OpenAI trained ChatGPT using Reinforcement Learning from Human Feedback (RLHF) chatgptintro2022 . The ChatGPT training process is shown in Figure 9 chatgptintro2022 . The original version of ChatGPT is fine-tuned from a model in the GPT-3.5 series, which is free and completed training in early 2022 chatgptintro2022 .
@@ 6.1 Performance of the models
Since ChatGPT is not only a classification model, we also asked the rationale when ChatGPT was identifying assumptions from the sentences. After analyzing the data, we can see that ChatGPT has made many errors in the classification task. Ten examples of incorrect classification with the rationale by ChatGPT are shown in Table 9. For E1, E2, E5, E7, and E9, the raw data qualify the SCA_IC1 criterion, which should be an SCA. In E1, although ChatGPT identified the keyword “assuming”, it misunderstood the sentence as a question. In E2, ChatGPT considered the sentence as “addressing the reader”, which is not true. In E5, the sentence includes “assume”, which is an assumption-related keyword, while ChatGPT said that the sentence does not use assumption-related keywords. In E7, ChatGPT explained the assumption as speculation and considered that it has to be confirmed. In E9, ChatGPT only saw part of the sentence: “something causes problems”. However, the sentence explicitly claimed an assumption that TensorFlow Probability’s JAX backend assumes properties like ‘numpy.int32 == tf.int32’ hold. For E3, ChatGPT identified an assumption-related keyword: “assumes” in the sentence, which does not exist. For E4 and E6, the sentence qualified the PA_IC1 criterion, while ChatGPT did not identify anything in the sentence. For E8 and E10, they are statements of facts, while ChatGPT considered them as assumptions. The reason is that sometimes ChatGPT would imagine the context of the sentence, even though we clearly asked it not to do that.

### 20 Mar 2024 v3

@@ 4.1 Research questions
Rationale: Algorithms usually have varying performance in classification problems, leading to different trained models. For example, Naive Bayes models are a group of probabilistic classifiers that have been widely used in text categorization since 1960. The Support Vector Machine algorithm is a commonly used supervised ML algorithm for data classification and regression analysis. To answer RQ1, we chose seven popular ML algorithms to train classification models for the identification of assumptions: Perceptron (Pct), Logistic Regression (LR), Linear Discriminant Analysis (LDA), K-Nearest Neighbors (KNN), Support Vector Machine (SVM), Naive Bayes (NB), and Classification and Regression Trees (CART). These algorithms have been used in similar studies, such as the classification of non-functional requirements Abad2017 and rationale Alkadhi2018 . Li et al. also used these seven ML algorithms to identify assumptions (binary classification) in the Hibernate project Li2019 . Compared to their work, the scale of the dataset used in this study is 19 times larger (15,384 vs. 800). Through this RQ, we plan to evaluate the performance of traditional ML models for identifying assumptions on the AssuEval dataset yangpackage2023 .
@@
Rationale: In addition to traditional ML and DL models, large language models (LLMs) are another potential solution for identifying assumptions. In this RQ, we chose one of the most representative LLMs (i.e., ChatGPT) to explore its performance in identifying assumptions. ChatGPT is an advanced AI chatbot developed by OpenAI [^3] https://openai.com/, which can interact in a conversational way chatgptintro2022 . As reported by OpenAI, more than 100 million people used ChatGPT. ChatGPT can be applied in various fields, such as answering specific questions, writing an essay, writing a song, writing code to implement a function, and code debugging chatgptintro2022 Meyer2023 . Specifically, we chose the free version [^4] https://chat.openai.com/chat of ChatGPT (based on GPT-3.5) as it is more available to the public. The results of this RQ show the performance of ChatGPT in identifying assumptions on the AssuEval dataset.
@@ 4.3 Data collection
The process of data collection is composed of five steps as described below.
(1) Select repositories. TensorFlow is one of the most popular DL frameworks, which is widely used in many DL systems and application domains. The TensorFlow project on GitHub started in 2015, having 161,680 commits [^5] https://github.com/tensorflow/tensorflow (accessed on 2024-3-19), 24,438 PRs [^6] https://github.com/tensorflow/tensorflow/pulls?q=is%3Apr (accessed on 2024-3-19), and 38,877 issues [^7] https://github.com/tensorflow/tensorflow/issues?q=is%3Aissue (accessed on 2024-3-19) till March 2024. Therefore, in this study, we chose the TensorFlow repository on GitHub as the main source to build our dataset. Moreover, Keras is an independent repository on GitHub, which is closely related to TensorFlow. For example, in some versions of TensorFlow, Keras has been integrated inside TensorFlow as a sub-module (i.e., tensorflow.keras). Therefore, in addition to TensorFlow, we also included the Keras repository on GitHub when constructing the AssuEval dataset yangpackage2023 .
(2) Determine the scope of the data. According to similar works (e.g., Zhang2018 Ruan2019 Han2020 ) that collect and analyze data on GitHub, in this study, we focus on commits, PRs, and issues of the selected repositories. There are also other sources (e.g., code comments) that may contain assumptions in the selected repositories. In this study, we do not aim to analyze assumptions (e.g., their importance and distribution) in DL framework projects. Instead, we aim to find a good algorithm and model to automatically identify assumptions in DL framework projects, and we consider that commits, PRs, and issues are representative data sources.
@@ 4.4.1 Construction process of the AssuEval dataset
(3) Conduct a pilot labeling. We conducted a pilot labeling on the TensorFlow and Keras repositories. The criteria were iteratively refined during the pilot study. For example, regarding an assumption that is a condition (e.g., “This obviously works if you can assume 0.0 is not real value in your data”), we decided not to label such data as SCA, because the assumption does not exist. Another example is that if a sentence is a question (e.g., “‘@willnorris for corporate CLAs, do we expect googlebot to say this has been signed? or do we just assume that if the user has @dropbox.com in their email that it is fine?”), which is about whether to make certain assumptions, we decided to label such data as NA, since the assumption does not exist. The third example is that if a sentence is a question (e.g., “@yaroslavvb I’m assuming that if I run the probe op in a session together with computation of a model this would return me the peek memory usage, is that correct?”), which is about the validity of an assumption rather than whether to make an assumption, we decided to label such data as SCA. More details of the included or excluded examples can be found in Table 3, Table 4, Table 5, and Table 6.
@@
(5) Label the data from the selected repositories. The first author manually labeled the data following the criteria using Assumption Miner. NAs are labeled as “0”, PAs are labeled as “1”, and SCAs are labeled as “2”.
@@ 4.4.2 Criteria of data labeling
E1: Cast image preprocessing inputs to compute dtype.
@@ 4.5.2 Model construction and training
For answering RQ1, we used the scikit-learn package to implement and train the seven ML models. The configurations of each model are shown in Table 7, which is based on the settings recommended in the scikit-learn package.
@@
To answer RQ2, we used the TensorFlow framework to train the ALBERT model. The configurations of the ALBERT model are shown in Table 8, which is based on the recommendation by the Google team on GitHub [^10] https://github.com/google-research/albert. The optimizer is Adam, the training batch size is 64, and the maximum sequence length of each sentence is 128. We trained the model to 10,000 epochs with an initial learning rate of 2e-5.
@@
To answer RQ3, we used the pre-trained ChatGPT model without retraining or further fine-tuning. OpenAI trained ChatGPT using Reinforcement Learning from Human Feedback (RLHF) chatgptintro2022 . The original version of ChatGPT is fine-tuned from a model in the GPT-3.5 series, which is free and completed training in early 2022 chatgptintro2022 .
@@ 4.6 Data analysis
Figure 7: TP, TN, FP, and FN of each label
In this study, we adopted a more strict standard to calculate classification performance, i.e., for TN, the predicted label for a sample should match exactly the corresponding label in the ground truth, as shown in Figure 8. “N/A” means “Not Applicable”. For instance, when examining the confusion matrix of NAs, if a sample is labeled SCA and the model classifies the sample as PA, even though it accurately identified negative labels (both SCA and PA are negative labels in this case), the model prediction is incorrect (PA) compared to the ground truth (SCA). Hence, it should not be considered a “True” negative.
@@ 6.1 Performance of the models
6.2 Reasons of low performance of ChatGPT
The ML models and the ALBERT model are more like a black box, which we do not know the rationale of how they identified the assumptions. Deferentially, ChatGPT is not only a classification model, but is also able to provide rationale behind its classification. Therefore, we also asked for the rationale when ChatGPT was identifying assumptions from sentences. After analyzing the data, we can see that ChatGPT has made many errors in the classification task. Ten examples of incorrect classification with the rationale of ChatGPT are shown in Table 10. For E1, E2, E5, E7, and E9, the raw data qualify the SCA_IC1 criterion, which should be an SCA. In E1, although ChatGPT identified the keyword “assuming”, it misunderstood the sentence as a question. In E2, ChatGPT considered the sentence as “addressing the reader”, which is not true. In E5, the sentence includes “assume”, which is an assumption-related keyword, while ChatGPT said that the sentence does not use assumption-related keywords. In E7, ChatGPT explained the assumption as speculation and considered that it has to be confirmed. In E9, ChatGPT only saw part of the sentence: “something causes problems”. However, the sentence explicitly claimed an assumption that TensorFlow Probability’s JAX backend assumes properties like ‘numpy.int32 == tf.int32’ hold. For E3, ChatGPT identified an assumption-related keyword: “assumes” in the sentence, which does not exist. For E4 and E6, the sentence qualified the PA_IC1 criterion, while ChatGPT did not identify anything in the sentence. For E8 and E10, they are statements of facts, while ChatGPT considered them as assumptions. The reason is that sometimes ChatGPT would imagine the context of the sentence, even though we clearly asked it not to do that.